Milan is a large city in Italy and it is now in a critical situation due to the increased number of cases with corona virus . this week 81% of the people in milan are "healthy" and, while 11% of the people are"sick" because they have got a positive test result ,and 8% represents the percentage of "death"cases . It is expected that 88% of the "healthy" people will stay "healthy" next week while 12% of the"healthy" people will be"sick" next week, it is expected also that in the next week , 19% of the "sick" people will be "healthy"(recovered cases),68% will stay"sick" and 13% will die. USING Markov analysis .find the percentage of the"death" cases for the next week .
Answer #1
From \ To Healthy Sick Dead
Healthy (0.81) 0.88 0.12 0.00
Sick (0.11) 0.19 0.68 0.13
Dead (0.08) 0.00 0.00 1.00

Prob{Healthy --> Dead} = 0.81*0 = 0

Prob{Sick --> Dead} = 0.11*0.13 = 0.0143

Prob{Dead --> Dead} = 0.08* 1 = 0.08

So, Prob{Dead in the next week} = 0 + 0.0143 + 0.08 = 0.943 or 9.43%